OpenStudy (bossimbacon):

What is the value of the expression when a = 5 and b = 3? 4a(2) - 5b (2) the () is exponents btw @DangerousJesse

OpenStudy (dangerousjesse):

Say `\(4\times a^{2}\)` To show \(4\times a^2\)

OpenStudy (bossimbacon):

oh.. kk @DangerousJesse

OpenStudy (dangerousjesse):

Plug in five and three \((4\times 5^2)-(5\times 3^2)\)

OpenStudy (dangerousjesse):

\(5\times 5=25\) and \(3\times 3=9\) \(5\times 25=100\) and \(5\times 9= 45\) Subtract \(100-45\)

OpenStudy (bossimbacon):

55!

OpenStudy (bossimbacon):

THANKS

OpenStudy (dangerousjesse):

Very good :) No problem.

OpenStudy (bossimbacon):

@DangerousJesse I'm done for today and i just finished my hw thanks to you! i will keep tagging you because you awesome! thanks again!